# Tractable and Provably Efficient Distributional Reinforcement Learning with General Value Function Approximation

### v1

@@ Abstract
Distributional reinforcement learning improves performance by effectively capturing environmental stochasticity, but a comprehensive theoretical understanding of its effectiveness remains elusive. In this paper, we present a regret analysis for distributional reinforcement learning with general value function approximation in a finite episodic Markov decision process setting. We first introduce a key notion of Bellman unbiasedness for a tractable and exactly learnable update via statistical functional dynamic programming. Our theoretical results show that approximating the infinite-dimensional return distribution with a finite number of moment functionals is the only method to learn the statistical information unbiasedly, including nonlinear statistical functionals. Second, we propose a provably efficient algorithm, SF-LSVI, achieving a regret bound of $\tilde{O}(d_{E}H^{\frac{3}{2}}\sqrt{K})$ where $H$ is the horizon, $K$ is the number of episodes, and $d_{E}$ is the eluder dimension of a function class.
1 Introduction
Distributional reinforcement learning (DistRL) [7, 41, 13, 30] is an advanced approach to reinforcement learning (RL) that focuses on the entire probability distribution of returns rather than solely on the expected return. By considering the full distribution of returns, distRL provides deeper insight into the uncertainty of each action, such as the mode or median. This framework enables us to make safer and more effective decisions that account for various risks [14, 46, 22, 29], particularly in complex real-world situations, such as robotic manipulation [10], neural response [36], stratospheric balloon navigation [8], algorithm discovery [19], and several game benchmarks [6, 34].
Although distRL has a solid theoretical basis and demonstrates effective empirical performance, its efficiency and tractability are still not fully understood. In practice, distributions contain an infinite amount of information, and we should resort to approximations using a finite number of parameters or statistical functionals, such as categorical [7] and quantile representations [16]. However, not all statistical functionals can be learned through the Bellman operator, as the meaning of statistical functionals is not typically preserved after updates. For example, the median is not preserved under the Bellman updates, as the median of a mixture of two distributions does not necessarily equate to the mixture of their medians. Hence, a distinct analysis is required to determine whether there exists a corresponding Bellman operator for each statistical functional that ensures commutativity. Rowland et al. [41] defined Bellman closedness as a property of statistical functionals that can be exactly learned by the existence of a corresponding Bellman operator.
At this point, we take a closer look at the distributional Bellman update and revisit what additional properties of statistical functionals beyond Bellman closedness are needed to construct tractable algorithms. In doing so, we identify and provide the following challenges that arise when statistical functionals are used to express the update process:
@@ 1 Introduction
Figure 1: Venn-Diagram of Statistical Functional Classes. The diagram shows the category of statistical functional. (Yellow) Within the linear statistical functional class, Rowland et al. [41] demonstrated that the solution of Bellman closedness is uniquely represented by polynomial exponential functionals. (Red) We extend this concept by introducing Bellman unbiasedness, which not only covers moment functionals but also central moment functionals from the broader class including nonlinear statistical functionals. (A) Maximum and minimum functionals are Bellman closed, while those are not unbiasedly estimatable. (B) Median and quantile functionals are neither Bellman closed and unbiased, highlighting that they are not proper to encode the distribution.
@@
  • Representing a mixture distribution with a finite, fixed number of parameters inevitably leads to approximation errors during the update process. For example, when expressing the mixture of two distributions, each represented by $N$ parameters, compressing $2N$ into $N$ parameters in the mixture results in inevitable information loss.
@@
  • Due to the unknown nature of the transition probability $\mathbb{P}(\cdot|s,a)$, the target distribution is estimated by sampling the next state $s^{\prime}$. For a tractable algorithm, it is crucial that the statistical functionals of the target distribution can be unbiasedly estimated using the statistical functionals from the sampled distribution.
@@
In this paper, we describe the statistical functional and sketch of the distribution, and introduce the additional approximation errors that can occur with the distRL method. We define the notion of Bellman unbiasedness, a property for representing the distribution tractably and prove that the moment functional is the only solution that satisfies this property, including nonlinear statistical functionals. Based on this property, we propose a provably efficient statistical functional reinforcement learning algorithm with general value function approximation, denoted as SF-LSVI. We will discuss the inherent issues of distributional Bellman completeness, a structural assumption previously defined in the literature [49, 11], and investigate the benefits of redesigning this concept using a sketch.
@@
  • Introduce a key property of Bellman unbiasedness for collections of statistical functionals (sketch) and show that the moment functionals are the unique structure that can be unbiasedly estimated and exactly learned from distributions, including nonlinear statistical functionals.
  • Revisit the existing concept of distributional Bellman Completeness through a statistical functional lens. We enhance the tractability by using moment functional, that the agent can only access a finite-dimensional embedding space. This removes the misspecification errors commonly encountered when discretizing infinite-dimensional distributions.
  • Propose a tractable and provably efficient distributional RL algorithm called SF-LSVI, achieving a tightest regret upper bound $\tilde{O}(d_{E}H^{\frac{3}{2}}\sqrt{K})$  [^1] We ignore poly-log terms in $H,S,A,K$ in the $\tilde{O}(\cdot)$ notation. in distributional framework. Our framework yields a better regret bound with a weaker structural assumption compared to prior results.
@@ 2 Related Work
Distributional RL.
In classical RL, the Bellman equation, which is based on expected returns, has a closed-form expression. However, it is questionable whether any statistical functionals of return distribution always have their corresponding closed-form expressions. Rowland et al. [41] introduced the notion of Bellman Closedness for collections of statistical functionals that can be updated in a closed form via Bellman update. They showed that the only Bellman-closed statistical functionals in the discounted setting are the moments $\mathbb{E}_{Z\sim\eta}[Z^{k}]$. More recently, Marthe et al. [35] proposed a general framework for distRL, where the agent plans to maximize their own utility functionals instead of expected return and formalized the property of Bellman Optimizability. They proved that the only $W_{1}$-continuous and linear Bellman Optimizable statistical functionals are shown to be exponential utilities $\frac{1}{\lambda}\log\mathbb{E}_{Z\sim\eta}[\exp(\lambda Z)]$ in the undiscounted setting.
In practice, C51 [7] and QR-DQN [16] are notable distributional RL algorithms where the convergence guarantees of sampled-based algorithms are proved [40, 42]. Dabney et al. [15] expanded the class of policies on arbitrary distortion risk measures by taking the based distribution non-uniformly and improve the sample efficiency from their implicit representation of the return distribution. Cho et al. [12] highlighted the drawbacks of optimistic exploration in distRL, introducing a randomized exploration method that perturbs the return distribution when the agent selects their next action.
Table 1: Comparison for different methods under distributional RL framework. $\mathcal{H}$ represents a subspace of infinite-dimensional space $\mathcal{F}^{\infty}$. To bound the eluder dimesion $d_{E}$, Wang et al. [49] and Chen et al. [11] assumed the discretized reward MDP.
@@
Tractability of
Distribution Estimation
@@
[11]
$\tilde{\mathcal{O}}(d_{E}H^{2}\sqrt{K})$  [^2] In Chen et al. [11], the regret bound is written as $\tilde{O}(d_{E}L_{\infty}(\rho)H\sqrt{K})$, where $L_{\infty}(\rho)$ represents the lipschitz constant of the risk measure $\rho$, i.e., $|\rho(Z)-\rho(Z^{\prime})|\leq L_{\infty}(\rho)\|F_{Z}-F_{Z^{\prime}}\|_{\infty}$. Since $L_{\infty}(\rho)\geq H$ in risk-neutral setting, we translate the regret bound into $\tilde{O}(d_{E}H^{2}\sqrt{K})$. $\text{dim}_{E}(\mathcal{H},\epsilon)$ distributional BC
@@
[3pt] SF-LSVI [Ours] $\tilde{\mathcal{O}}(d_{E}H^{\frac{3}{2}}\sqrt{K})$ $\text{dim}_{E}(\mathcal{F}^{N},\epsilon)$ statistical functional BC none
@@
Regret bounds have been studied for a long time in RL, across various domains such as bandit [32, 1, 43], tabular RL [28, 3, 38, 39, 26], and linear function approximation [27, 51, 52]. In recent years, deep RL has shown significant performance using deep neural networks as function approximators, and attempts have been made to analyze whether it is efficient in terms of general function approximation. Wang et al. [50] established a provably efficient RL algorithm with general value function approximation based on the eluder dimension $d_{E}$ [43] and achieves a regret upper bound of $\tilde{O}(\text{poly}(d_{E}H)\sqrt{K})$. To circumvent the intractability from computing the upper confidence bound, Ishfaq et al. [25] injected the stochasticity on the training data and get the optimistic value function instead of upper confidence bound, enhancing computationally efficiency. Beyond risk-neutral setting, several prior works have shown regret bounds under risk-sensitive objectives (e.g., entropic risk [20, 33], CVaR [5]), which align with our approach in that they are built on a distribution framework. Liang and Luo [33] achieved the regret upper bound of $\tilde{O}(\exp(H)\sqrt{|\mathcal{S}|^{2}|\mathcal{A}|H^{2}K})$ and the lower bound of $\Omega(\exp(H)\sqrt{|\mathcal{S}||\mathcal{A}|HK})$ in tabular setting.
@@
Recently, only few efforts have aimed to bridge the gap between two fields. Wang et al. [49] proposed a distributional RL algorithm, O-DISCO, which enjoys small-loss bound by using a log-likelihood objective. Similarly, Chen et al. [11] provided a risk-sensitive reinforcement learning framework with static lipschitz risk measure. While these studies analyze within a distributional framework, they do not address the intractability of implementation in infinite-dimensional space of distributions. In contrast, our approach focuses on a statistical functional framework, providing a detailed comparison with other distRL methods as shown in Table 1.
@@ 3 Preliminaries
$Z_{h}^{\pi}(s,a)(:\mathcal{S}\times\mathcal{A}\times\Omega\rightarrow\mathbb{R})\coloneqq\sum_{h^{\prime}=h}^{H}r_{h^{\prime}}(s_{h^{\prime}},a_{h^{\prime}})\mid s_{h}=s,a_{h}=a,a_{h^{\prime}}=\pi_{h^{\prime}}(s_{h^{\prime}}),$
@@
$\bar{Z}_{h}^{\pi}(s)(:\mathcal{S}\times\Omega\rightarrow\mathbb{R})\coloneqq\sum_{h^{\prime}=h}^{H}r_{h^{\prime}}(s_{h^{\prime}},a_{h^{\prime}})\mid s_{h}=s,a_{h^{\prime}}=\pi_{h^{\prime}}(s_{h^{\prime}}),$
@@
Note that $\bar{Z}^{\pi}_{h}(s)=Z^{\pi}_{h}(s,\pi(s))$ and $\bar{\eta}_{h}^{\pi}(s)=\eta_{h}^{\pi}(s,\pi(s))$. We use $\pi^{\star}$ to denote an optimal policy and denote $V_{h}^{\star}(s)=V_{h}^{\pi^{\star}}(s),Q_{h}^{\star}(s,a)=Q_{h}^{\pi^{\star}}(s,a),\eta_{h}^{\star}(s,a)=\eta_{h}^{\pi^{\star}}(s,a)$, and $\bar{\eta}_{h}^{\star}(s)=\bar{\eta}_{h}^{\pi^{\star}}(s)$. For notational simplicity, we denote the expectation over transition, $[\mathbb{P}_{h}V_{h+1}^{\pi}](s,a)=\mathbb{E}_{s^{\prime}\sim\mathbb{P}_{h}(\cdot|s,a)}V_{h+1}^{\pi}(s^{\prime})$, $[\mathbb{P}_{h}\bar{Z}_{h+1}^{\pi}](s,a)=\mathbb{E}_{s^{\prime}\sim\mathbb{P}_{h}(\cdot|s,a)}\bar{Z}_{h+1}^{\pi}(s^{\prime})$, and $[\mathbb{P}_{h}\bar{\eta}_{h+1}^{\pi}](s,a)=\mathbb{E}_{s^{\prime}\sim\mathbb{P}_{h}(\cdot|s,a)}\bar{\eta}_{h+1}^{\pi}(s^{\prime}).$  [^3] Note that $\mathbb{E}_{s^{\prime}\sim\mathbb{P}_{h}(\cdot|s,a)}\bar{\eta}^{\pi}_{h+1}(s^{\prime})$ is a mixture distribution. For brevity, we refer to $\bar{\eta}^{\pi}$ simply as $\bar{\eta}$.
@@
Recall that $\eta^{\star}_{h}$ satisfies the distributional Bellman optimality equation:
$\eta_{h}^{\star}(s,a)=(\mathcal{T}_{h}\eta_{h+1}^{\star})(s,a)$ $=\mathbb{E}_{s^{\prime}\sim\mathbb{P}_{h}(\cdot|s,a),a^{\prime}\sim\pi_{h}^{\star}(\cdot|s^{\prime})}[(\mathcal{B}_{r_{h}})_{\#}\eta_{h+1}^{\star}(s^{\prime},a^{\prime})]$
@@
where $\mathcal{B}_{r}:\mathbb{R}\rightarrow\mathbb{R}$ is defined by $\mathcal{B}_{r}(x)=r+x$, and $g_{\#}\eta\in\mathscr{P}(\mathbb{R})$ is the pushforward of the distribution $\eta$ through $g$, i.e., $g_{\#}\eta(A)=\eta(g^{-1}(A))$ for any Borel set $A\subseteq\mathbb{R}$.
@@
For a given $N$, we denote an $N-$dimensional function class $\mathcal{F}^{N}\coloneqq\mathcal{F}^{(1)}\times\dots\times\mathcal{F}^{(N)}\subseteq\Big{\{}f=[f^{(1)},\cdots,f^{(N)}]:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}^{N}\Big{\}}.$ Given a dataset $\mathcal{D}=\{(s_{t},a_{t},[z_{t}^{(1)},\dots,z_{t}^{(N)}])\}_{t=1}^{|\mathcal{D}|}\subseteq\mathcal{S}\times\mathcal{A}\times\mathbb{R}^{N}$, a set of state-action pairs $\mathcal{Z}=\{(s_{t},a_{t})\}_{t=1}^{|\mathcal{Z}|}\subseteq\mathcal{S}\times\mathcal{A}$ and for a function $f:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}^{N}$, we define the norm $\|f^{(n)}\|_{\infty},\|f\|_{\infty,1},\|f\|_{\mathcal{D}},\|f\|_{\mathcal{Z}}$ as written in Appendix A. For a set of (vector-valued) functions $\mathcal{F}^{N}\subseteq\{f:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}^{N}\}$, the width function of $(s,a)$ is defined as $w^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(n)}}(\mathcal{F}^{N},s,a)\coloneqq\max_{f,g\in\mathcal{F}^{N}}|f^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(n)}}(s,a)-g^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(n)}}(s,a)|$.

### v2

# Bellman Unbiasedness: Toward Provably Efficient Distributional Reinforcement Learning with General Value Function Approximation

@@ Abstract
Distributional reinforcement learning improves performance by capturing environmental stochasticity, but a comprehensive theoretical understanding of its effectiveness remains elusive. In addition, the intractable element of the infinite dimensionality of distributions has been overlooked. In this paper, we present a regret analysis of distributional reinforcement learning with general value function approximation in a finite episodic Markov decision process setting. We first introduce a key notion of Bellman unbiasedness which is essential for exactly learnable and provably efficient distributional updates in an online manner. Among all types of statistical functionals for representing infinite-dimensional return distributions, our theoretical results demonstrate that only moment functionals can exactly capture the statistical information. Secondly, we propose a provably efficient algorithm, SF-LSVI, that achieves a tight regret bound of $\tilde{O}(d_{E}H^{\frac{3}{2}}\sqrt{K})$ where $H$ is the horizon, $K$ is the number of episodes, and $d_{E}$ is the eluder dimension of a function class.
Machine Learning, ICML
@@
Figure 1: Venn-Diagram of Statistical Functional Classes. The diagram illustrates categories of statistical functional. (Yellow $\cap$ Blue) Within the linear statistical functional class, Rowland et al. (2019) showed that the only functionals satisfying Bellman closedness are moment functionals. (Red $\cap$ Blue) We extend this concept by introducing the notion of Bellman unbiasedness, which not only covers moment functionals but also includes central moment functionals from the broader class including nonlinear statistical functionals. (Yellow $\cap$ $\text{Blue}^{c}$) According to Lemmas 3.2 and 4.4 of Rowland et al. (2019), categorical functionals are linear but not Bellman closed. (A) Maximum and minimum functionals are Bellman closed, while they are not unbiasedly estimatable. (B) Median and quantile functionals are neither Bellman closed nor unbiased, highlighting that they are not proper to encode the distribution in terms of exactness. The proofs corresponding to each region are provided in Appendix C.
Distributional reinforcement learning (DistRL) (Bellemare et al., 2017; Rowland et al., 2019; Choi et al., 2019; Kim et al., 2024) is an advanced approach to reinforcement learning (RL) that focuses on the entire probability distribution of returns rather than solely on the expected return. By considering the full distribution of returns, distRL provides deeper insight into the uncertainty of each action, such as the mode or median. This framework enables us to make safer and more effective decisions that account for various risks (Chow et al., 2015; Son et al., 2021; Greenberg et al., 2022; Kim et al., 2023), particularly in complex real-world situations, such as robotic manipulation (Bodnar et al., 2019), neural response (Muller et al., 2024), stratospheric balloon navigation (Bellemare et al., 2020), algorithm discovery (Fawzi et al., 2022), and several game benchmarks (Bellemare et al., 2013; Machado et al., 2018). While the distributional approach offers richer information, two key theoretical challenges are introduced that distinguish it from expectation-based RL.
Infinite-dimensionality of distribution.
In practice, distributions contain an infinite amount of information, and we must resort to approximations using a finite number of parameters or statistical functionals, such as categorical (Bellemare et al., 2017) and quantile representations (Dabney et al., 2018b). However, previous works often conducted analyses while overlooking these intractable nature of distributions. Additionally, not all statistical functionals can be exactly learned through the Bellman operator, as the meaning of statistical functionals is not preserved after updates. For example, the median is not preserved under the Bellman updates, as the median of a mixture of two distributions does not equal the mixture of their medians. Thus, a fundamental question arises: For a given statistical functional, does there exist a corresponding Bellman operator that ensures exactness? To formalize this issue, Rowland et al. (2019) introduced Bellman closedness, which characterizes statistical functionals that can be exactly learned in the presence of a corresponding Bellman operator.
Online distributional update.
In this paper, we focus on developing an algorithm that efficiently explores from a regret minimization perspective while simultaneously performing distributional Bellman updates in an online manner. One possible approach to addressing this problem is to first update the policy using an existing provably efficient non-distributional RL algorithm and then estimate the distribution via additional rollouts. However, decoupling these two processes introduces several drawbacks. First, adding extra rollouts solely for distribution estimation is sample-inefficient, and the limited number of rollouts inevitably introduce accumulated approximation errors in the estimated distribution throughout the learning process. Moreover, the estimation is confined to the return distribution of the executed policy, making it difficult to reuse for estimating distributions under different policies, thereby moving further away from off-policyness.
To overcome those two fundamental challenges inherent to DistRL, we take a closer look at the distributional Bellman update and revisit what additional properties of statistical functionals, beyond Bellman closedness, are required to construct online DistRL algorithms that are not only exactly learnable but also provably efficient in terms of regret. In this context, we identify the following additional issues that arise when using statistical functionals for updates instead of the full distribution:
@@
  • Representing a mixture distribution with a finite, fixed number of parameters leads to approximation errors during the update. For example, when expressing the mixture of two distributions, each represented by $N$ parameters, compressing $2N$ into $N$ parameters in the mixture results in inevitable information loss.
@@
  • Due to the unknown nature of the transition $\mathbb{P}(\cdot|s,a)$, the target distribution is estimated by sampling the next state $s^{\prime}$. Hence, the statistical functionals of the target distribution should be unbiasedly estimated using the statistical functionals from the sampled distribution.
@@
In this paper, we introduce a key concept, Bellman unbiasedness, for precise information learnability of a distribution from a finite number of samples in an online setting. As shown in Figure 1, we prove that the moment functional remains the only solution in a class that includes nonlinear statistical functionals that satisfies both properties. We then discuss the inherent intractability of distributional Bellman completeness (distBC) – a structural assumption previously defined in the literature (Wang et al., 2023; Chen et al., 2024) – and investigate the benefits of redesigning this concept using a collection of statistical functionals. Finally, we propose a provably efficient statistical functional RL algorithm with general value function approximation, called SF-LSVI.
@@
  • Introduce a key property of Bellman unbiasedness for exactly learnable and provably efficient online distRL algorithm. We show that the moment functional is the unique structure in a class including nonlinear statistical functionals.
  • Describe the inherent intractability of infinite-dimensional distributions and analyze how hidden approximation error prevents the design of provably efficient algorithms. To address this, we revisit the existing structural assumption of distributional Bellman Completeness through a statistical functional lens.
  • Propose exactly learnable and provably efficient distRL algorithm called SF-LSVI, achieving a tight regret upper bound $\tilde{O}(d_{E}H^{\frac{3}{2}}\sqrt{K})$.  [^1] We ignore poly-log terms in $H,S,A,K$ in the $\tilde{O}(\cdot)$ notation. Our framework yields a tighter regret bound with a weaker structural assumption compared to prior results in distRL.
@@ 1 Related Work
Table 1: Comparison for different methods under distributional RL framework. $\mathcal{H}$ represents a subspace of infinite-dimensional space $\mathcal{F}^{\infty}$. To bound the eluder dimesion $d_{E}$, Wang et al. (2023) and Chen et al. (2024) assumed the discretized reward MDP.
@@
Finite
Representation
Exactly
Learnable
@@
(Chen et al., 2024)
$\tilde{\mathcal{O}}(d_{E}H^{2}\sqrt{K})$  [^2] In Chen et al. (2024), the regret bound is written as $\tilde{O}(d_{E}L_{\infty}(\rho)H\sqrt{K})$, where $L_{\infty}(\rho)$ represents the lipschitz constant of the risk measure $\rho$, i.e., $|\rho(Z)-\rho(Z^{\prime})|\leq L_{\infty}(\rho)\|F_{Z}-F_{Z^{\prime}}\|_{\infty}$. Since $L_{\infty}(\rho)\geq H$ in risk-neutral setting, we translate the regret bound into $\tilde{O}(d_{E}H^{2}\sqrt{K})$. $\text{dim}_{E}(\mathcal{H},\epsilon)$ distributional BC
@@
SF-LSVI
[Ours]
$\tilde{\mathcal{O}}(d_{E}H^{\frac{3}{2}}\sqrt{K})$ $\text{dim}_{E}(\mathcal{F}^{N},\epsilon)$ statistical functional BC none
Distributional RL.
In classical RL, the Bellman equation, which is based on expected returns, has a closed-form expression. However, it remains unclear whether any statistical functionals of return distribution always have their corresponding closed-form expressions. Rowland et al. (2019) introduced the notion of Bellman closedness for collections of statistical functionals that can be updated in a closed form via Bellman update. They showed that the only Bellman-closed statistical functionals in the discounted setting are the moments $\mathbb{E}_{Z\sim\eta}[Z^{k}]$. More recently, Marthe et al. (2023) proposed a general framework for distRL, where the agent plans to maximize its own utility functionals instead of expected return, formalizing this property as Bellman Optimizability. They further demonstrated that in the undiscounted setting, the only $W_{1}$-continuous and linear Bellman optimizable statistical functionals are exponential utilities $\frac{1}{\lambda}\log\mathbb{E}_{Z\sim\eta}[\exp(\lambda Z)]$.
In practice, C51 (Bellemare et al., 2017) and QR-DQN (Dabney et al., 2018b) are notable distributional RL algorithms where the convergence guarantees of sampled-based algorithms are proved (Rowland et al., 2018, 2023). Dabney et al. (2018a) expanded the class of policies on arbitrary distortion risk measures by taking the based distribution non-uniformly and improve the sample efficiency from their implicit representation of the return distribution. Cho et al. (2023) highlighted the drawbacks of optimistic exploration in distRL, introducing a randomized exploration that perturbs the return distribution when the agent selects their next action.
@@
Regret bounds have been studied for a long time in online RL, across various domains such as bandit (Lattimore & Szepesvári, 2020; Abbasi-Yadkori et al., 2011; Russo & Van Roy, 2013), tabular RL (Kakade, 2003; Auer et al., 2008; Osband & Van Roy, 2016; Osband et al., 2019; Jin et al., 2018), and linear function approximation (Jin et al., 2020; Wang et al., 2019; Zanette et al., 2020). In recent years, deep RL has shown significant performance using deep neural networks as function approximators, and attempts have been made to analyze whether it is efficient in terms of general function approximation (Jin et al., 2021; Agarwal et al., 2023). Wang et al. (2020) established a provably efficient RL algorithm with general value function approximation based on the eluder dimension $d_{E}$ (Russo & Van Roy, 2013) and achieves a regret upper bound of $\tilde{O}(\text{poly}(d_{E}H)\sqrt{K})$. To circumvent the intractability from computing the upper confidence bound, Ishfaq et al. (2021) injected the stochasticity on the training data and get the optimistic value function instead of upper confidence bound, enhancing computationally efficiency. Beyond risk-neutral setting, several prior works have shown regret bounds under risk-sensitive objectives (e.g., entropic risk (Fei et al., 2021; Liang & Luo, 2022), CVaR (Bastani et al., 2022)), which align with our approach in that they are built on a distribution framework. Liang & Luo (2022) achieved the regret upper bound of $\tilde{O}(\exp(H)\sqrt{|\mathcal{S}|^{2}|\mathcal{A}|H^{2}K})$ and the lower bound of $\Omega(\exp(H)\sqrt{|\mathcal{S}||\mathcal{A}|HK})$ in tabular setting.
@@
Recently, only few efforts have aimed to bridge the gap between two fields. Wang et al. (2023) proposed a distributional RL algorithm, O-DISCO, which enjoys small-loss bound by using a log-likelihood objective. Similarly, Chen et al. (2024) provided a risk-sensitive RL framework with static lipschitz risk measure. While these studies analyze within a distributional framework, they do not address the intractability of implementation in infinite-dimensional space of distributions. In contrast, our approach focuses on a statistical functional framework, providing a detailed comparison with other distRL methods as shown in Table 1.
@@ 2 Preliminaries
$Z_{h}^{\pi}(s,a)(:\mathcal{S}\times\mathcal{A}\times\Omega\rightarrow\mathbb{R})$
$\coloneqq\sum_{h^{\prime}=h}^{H}r_{h^{\prime}}(s_{h^{\prime}},a_{h^{\prime}})\mid s_{h}=s,a_{h}=a,a_{h^{\prime}}=\pi_{h^{\prime}}(s_{h^{\prime}}),$
@@
$\bar{Z}_{h}^{\pi}(s)(:\mathcal{S}\times\Omega\rightarrow\mathbb{R})$
$\coloneqq\sum_{h^{\prime}=h}^{H}r_{h^{\prime}}(s_{h^{\prime}},a_{h^{\prime}})\mid s_{h}=s,a_{h^{\prime}}=\pi_{h^{\prime}}(s_{h^{\prime}}),$
@@
Note that $\bar{Z}^{\pi}_{h}(s)=Z^{\pi}_{h}(s,\pi(s))$ and $\bar{\eta}_{h}^{\pi}(s)=\eta_{h}^{\pi}(s,\pi(s))$. We use $\pi^{\star}$ to denote an optimal policy ( i.e., $){\pi^{\star}_{h}}(\cdot|s)=\arg\max_{\pi}V^{\pi}_{h}(s)$ and denote $V_{h}^{\star}(s)=V_{h}^{\pi^{\star}}(s),Q_{h}^{\star}(s,a)=Q_{h}^{\pi^{\star}}(s,a)$, $\eta_{h}^{\star}(s,a)=\eta_{h}^{\pi^{\star}}(s,a)$, and $\bar{\eta}_{h}^{\star}(s)=\bar{\eta}_{h}^{\pi^{\star}}(s)$. For notational simplicity, we denote the expectation over transition, $[\mathbb{P}_{h}V_{h+1}^{\pi}](s,a)=\mathbb{E}_{s^{\prime}\sim\mathbb{P}_{h}(\cdot|s,a)}V_{h+1}^{\pi}(s^{\prime})$, $[\mathbb{P}_{h}\bar{Z}_{h+1}^{\pi}](s,a)=\mathbb{E}_{s^{\prime}\sim\mathbb{P}_{h}(\cdot|s,a)}\bar{Z}_{h+1}^{\pi}(s^{\prime})$, and $[\mathbb{P}_{h}\bar{\eta}_{h+1}^{\pi}](s,a)=\mathbb{E}_{s^{\prime}\sim\mathbb{P}_{h}(\cdot|s,a)}\bar{\eta}_{h+1}^{\pi}(s^{\prime}).$  [^3] Note that $\mathbb{E}_{s^{\prime}\sim\mathbb{P}_{h}(\cdot|s,a)}\bar{\eta}^{\pi}_{h+1}(s^{\prime})$ is a mixture distribution. For brevity, we refer to $\bar{\eta}^{\pi}$ simply as $\bar{\eta}$.
@@
Recall that $\eta^{\star}_{h}$ satisfies the following optimality equation:
$\eta_{h}^{\star}(s,a)$ $=(\mathcal{T}_{h}\eta_{h+1}^{\star})(s,a)$
$\coloneqq\mathbb{E}_{s^{\prime}\sim\mathbb{P}_{h}(\cdot|s,a),a^{\prime}\sim\pi_{h}^{\star}(\cdot|s^{\prime})}[(\mathcal{B}_{r_{h}})_{\#}\eta_{h+1}^{\star}(s^{\prime},a^{\prime})]$
@@
where $\mathcal{B}_{r}:\mathbb{R}\rightarrow\mathbb{R}$ is defined by $\mathcal{B}_{r}(x)=r+x$, and $g_{\#}\eta\in\mathscr{P}(\mathbb{R})$ is the pushforward of the distribution $\eta$ through $g$ (i.e., $g_{\#}\eta(A)=\eta(g^{-1}(A))$ for any Borel set $A\subseteq\mathbb{R}$).
@@
For a given $N$, we denote an $N-$dimensional function class $\mathcal{F}^{N}\coloneqq\mathcal{F}^{(1)}\times\dots\times\mathcal{F}^{(N)}\subseteq\Big{\{}f=[f^{(1)},\cdots,f^{(N)}]:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}^{N}\Big{\}}.$ Given a dataset $\mathcal{D}=\{(s_{t},a_{t},[z_{t}^{(1)},\dots,z_{t}^{(N)}])\}_{t=1}^{|\mathcal{D}|}\subseteq\mathcal{S}\times\mathcal{A}\times\mathbb{R}^{N}$, a set of state-action pairs $\mathcal{Z}=\{(s_{t},a_{t})\}_{t=1}^{|\mathcal{Z}|}\subseteq\mathcal{S}\times\mathcal{A}$ and for a function $f:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}^{N}$, we define the norm $\|f^{(n)}\|_{\infty},\|f\|_{\infty,1},\|f\|_{\mathcal{D}},\|f\|_{\mathcal{Z}}$ as written in Appendix A. For a set of (vector-valued) functions $\mathcal{F}^{N}\subseteq\{f:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}^{N}\}$, the width function of $(s,a)$ is defined as $w^{{(n)}}(\mathcal{F}^{N},s,a)\coloneqq\max_{f,g\in\mathcal{F}^{N}}|f^{{(n)}}(s,a)-g^{{(n)}}(s,a)|$.